YourlocalRandom2000:

Help what is [\int_{0}^{\pi} \sin(x) \cos(x) ,dx]

Makaylachuck23backup:

oh thats

Makaylachuck23backup:

JRUEREUERE828827262629020332423234=000=0+++//./...\.\.

5StarFab:

Well as we can see it's letter's, numbers and symbols

5StarFab:

letters'

OLIVER69:

\[[\int\limits_{0}^{\pi} \sin(x) \cos(x) ,dx]\] this is what that is and I have ni idea how to solve that

starla:

this is app to solve math

surjithayer:

\[let~I=\int\limits_{0}^{\pi} \sin x \cos x~dx\] \[=\frac{ 1 }{ 2 }\int\limits_{0}^{\pi}2\sin x \cos x ~dx\] \[=\frac{ 1 }{ 2 }\int\limits_{0}^{\pi}\sin 2x dx\] put 2x =t 2dx=dt when x=0,t=0 when x=pi t=2 pi \[I=\frac{ 1 }{ 2 }\int\limits_{0}^{2\pi}\sin t~\frac{ dt }{ 2 }=\frac{ 1 }{ 4}\int\limits_{0}^{2 \pi} \sin t~dt\] \[\sin (2\pi-t)=\sin t \] \[I=2\times \frac{ 1 }{ 4}\int\limits_{0}^{\pi}\sin t dt=\frac{ 1 }{ 2 }\int\limits_{0}^{\pi}\sin t~dt\] \[\sin (\pi-t)=\sin t\] \[I=2\times \frac{ 1 }{2 }\int\limits_{0}^{\frac{ \pi }{2 }}\sin t~dt=[-\cos t] ~0~\to~\frac{ \pi }{ 2}\] \[=-(\cos \frac{ \pi }{ 2 }-\cos 0)=-(0-1)=1\]